OpenStudy (anonymous):

A periscope is 5 feet above the surface of the ocean. Through it can be seen a ship that rises to 50 feet above the water. To the nearest mile, the farthest away that the ship could be is ( ) miles

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

5^2+50^2=25+2500=2525 2525−−−−√ =?

OpenStudy (anonymous):

50.24?

OpenStudy (anonymous):

No not quite.

OpenStudy (anonymous):

Try to work it out. :]

OpenStudy (anonymous):

How to work it out?

OpenStudy (anonymous):

I'll give you a hint; it's in the 500's, you should be able to get it from there.

OpenStudy (anonymous):

505

OpenStudy (anonymous):

Am I correct? @Aimee98

OpenStudy (anonymous):

Yup you're right!! Great job! :D Here's a medal.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

505 is incorrect

OpenStudy (anonymous):

@cwrw238